OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1) explain the process for solving this equation. Please include your work at each step of the process along with the final answer

OpenStudy (anonymous):

alright with this problem you first start with 2(x+6). You wanna mulitiply 2 and x. then 2 and 6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok the problem should now look like this: 2x+6-10=12+4(x-1) Does your problem look like this

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

then you go to the 4(x-1) part and multiply 4 and the x, and mulitply 4 and the 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that the final answer

OpenStudy (anonymous):

no there more steps.

OpenStudy (anonymous):

next you should have 2x+6-10=12+4x-4

OpenStudy (anonymous):

x=-3 RUINED THE STORY!

OpenStudy (anonymous):

no x=-6

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

no its -3

OpenStudy (anonymous):

99% sure id check your math

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

thanx 4 helping ^_^

OpenStudy (anonymous):

ok im sure its -3

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i gues you were rite

OpenStudy (anonymous):

i think you messed up the left side somewhere in there

OpenStudy (anonymous):

so whats the answer now :c

OpenStudy (anonymous):

its -3

OpenStudy (anonymous):

is it the same steps though

OpenStudy (anonymous):

lemme type them out 2 seconds

OpenStudy (anonymous):

2x+12-10=12+4x-4 2x+2=8+4x -2x=6 x=-3